OpenStudy (anonymous):

Which box-and-whisker plot best represents the data? 34, 36, 43, 46, 47, 48, 48, 52, 53, 54 A. VHS_PA_S2_07_L207_L306_LQ_Q2a.gif B. Box-and-Whisker Plots C. Box-and-Whisker Plots D. Box-and-Whisker Plots

OpenStudy (texaschic101):

34,36,43, 46, 47, 48, 48, 52,53,54 Q2, or the median, is (47 + 48)/2 = 95/2 = 47.5 Q1 = 43 Q3 = 52 maximum = 54 minimum = 34 it should look like this..